Recent Log Entries

devfs RAID card replaced - back online

Log Number: 3859003Logbook(s): Entry Makers: cuffeadamcBody: 

devfs is back online with hopefully no more outages.

Details:
After the second round of bus resets, I replaced the RAID card in devfs that serves up /apps and re-seated the PCIe riser. This replacement went well and the system shows no more signs of issues. Will monitor and cut--over to the new devfs as the next step. I wanted to avoid this as a first step since it would require rebooting all client systems. With the repair, they should recover without needing a reboot.

INJ UHV Readbacks Unresolved

Log Number: 3859001Logbook(s): Body: 

Readbacks for the UHV Ion Pump Supplies were unresolved. The supplies use the serial connection of the transition module for iocin1. It looks like the ioc was offline. After rebooting iocin1, all UHV signals are now resolving.

devfs outage continuing - failures returned

Log Number: 3859000Logbook(s): Body: 

The issues with devfs have returned. Replacing the RAID card now. This will hopefully resolve the issues if not, we will migrate to the new fileserver. System should be back online within 30 minutes.

devfs back online - monitoring

Log Number: 3858999Logbook(s): Entry Makers: cuffeadamcBody: 

devfs has a PCIe bus error this morning (12/02/2020) at 03:03. This slot with the error contains the RAID card that is home to the /apps (pubtools,cuse,etc...) file systems. This caused the OS to set the file system as write-locked and was causing I/O errors. The system is back on-line and we are monitoring it.

-------------------------------------------------------------------------------
Recovery Operations:
-------------------------------------------------------------------------------
System harware log was cleared:
racadm clrsel

Investigating devfs issues

Log Number: 3858995Logbook(s): Entry Makers: cuffeBody: 

The dev fiefdom file server, devfs, is experiencing storage related issues. We are investigating and will likely need to reboot the server. A reboot would result in a temporary service outage for the dev fiefdom clients and users with home directories in devuser.

Someone will post follow up entries as more is known.

ABIL Tag #4509 added to Macropule chassis.

Log Number: 3858994Logbook(s): Body: The following entry was placed in the Accelerator Bypassed-Interlocks Log (ABIL) Device: Macropule chassis Function: Bypass laser interlocks Location: Injector Purpose: allow laser operation Exp.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858992Logbook(s): Body: 

Day #51, Week 11 Day 3, Plans:

Tunnel --
* SSG and CIS working together to enable use of CEBAF
laser for Gun->Y-chamber alignment
* CIS installing laser vacuum window to Y-chamber
*

InjSB --
* Gun HV interlock work, IN01B{02-05} (Windham)
* PSS Upgrades, IN02B{NI,29,30} +briefly elsewhere,
afternoons (Beizer)

TL --
* V-wien HV testing at UITF continues, flipping polarities
today, then back up to+/-20kV (Hernandez-Garcia)
* New prebuncher with RF group for testing/tuning
(Wissman)
* A1/A2 cube assembly (Adderley)

Toolbox attended by: Hofler, Hernandez-Garcia, Gould,
Norris, Croke, Col

Beam Authorization: Credited Control Downgraded

Log Number: 3858986Logbook(s): Tags: Body: Credited Control: Permanent Shielding Beam Destinations: UITF KeV Dump (IDLK501)UITF MeV (IDLM601) [Waist Height]UITF MeV (IDLM703) [Waist Height]UITF HDIce (IDLMB02) [Elevated] Modified On: 02-Dec-2020 08:33 Modified By: Welch, Keith (welch) Verification: Not Verified Comments: Roof tiles removed. See: Beam Authorization

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858969Logbook(s): Body: 

Day #50, Week 11 Day 2, Wrap-Up:

AIPINJ --
* New Y-chamber and solenoid installed in tunnel by
MechInstall. Pictures -- 3858956
* Gun table moved onto new base pads
* V-wien HV testing at UITF going well -- with electrodes at
+/- 20kV ~1% rise in vacuum compared to background.
Electrodes left soaking overnight. Test will be repeated
tomorrow, flipping electrode polarities. 3858966
* H-wien end plates that were found to undesirably leak
magnetic field stowed away in TL. New Nickel 200 end
plates ordered, enough for H-wien and spares.

V-Wien HV tested to +/-20kV @ UITF

Log Number: 3858966Logbook(s): Entry Makers: chgarciaBody: 

V-Wien high voltage testing started today. With 10% overhead voltage per electrode (+/- 20kV DC) of that required for 90 degree spin rotation, the vacuum level change was about only 1% compared to background. The electrodes will be left soaking overnight at these voltages. Wednesday morning, the voltage will be ramped down, the polarity switched, and the test repeated.

The test started by applying -20kV to the right plate (per beam direction convention) at 250 volts/minute ramp, while the opposite plate was kept at 0 volts.

COMMENT ON *Vacuum Activities CEBAF 03/03/20 through

Log Number: 3858960Logbook(s): Body: 

1L02 final insulating vacuum leak check good. Minimal detectable leak 1e-9. Sprayed all waveguide top hat assemblies/rotary feedthroughs, access ports o-rings, upstream/downstream beampipe o-rings, end can assembly/o-ring.

1L03 Installed ceramic warm windows on cavities 1-8.

AIPINJ region 1 in the tunnel

Log Number: 3858956Logbook(s): Entry Makers: gramesBody: 

A major milestone....the first few meters (region 1) of the injector upgrade is now in the tunnel.

Great job to all!

Pages